What is the sum of the infinite geometric series?
(-15)+ 0+ 15 +30+.....+195
Please show me step by step, thank you in advance!

Answer by richwmiller(17219)

This is not a geometric series.
An = a1 + (n - 1)d
195 = -15 + (14)15
and it has 15 terms
the difference is 15
195 is the 15th term starting with -15
It is a finite arithmetic series
S = ½(a1 + an)n
S = ½(-15 + 195)15
S = 90*15=1350